One of two supplementary angles is 8 more than the other. Find the measure of the larger angle.
Answer by nycsub_teacher(90)

One of two supplementary angles is 8 more than the other. Find the measure of the larger angle.
==========================================================
Since the angles are supplementary, then they add up to 180.
Let x = smaller angle
Let x + 8 = larger angle
We now have this:
x + x + 8 = 180
2x + 8 = 180
2x = 180 - 8
2x = 172
x = 172/2
x = 86
The larger angle is x + 8 and I just x to be 86.
So, the larger angle is 86 + 8 = 94